two planes leave from Dallas, Texas. The first plane travels east at a rate of 250 mph. The second plane travels west at a speed of 225 mph. How long will it take for them to be 2850 miles apart?

Answer by stanbon(57346)

two planes leave from Dallas, Texas. The first plane travels east at a rate of 250 mph. The second plane travels west at a speed of 225 mph. How long will it take for them to be 2850 miles apart?
---------------
They are separating at the rate of 250+225 = 475 mph
---
time = distance/rate
---
time = 2850/475 = 6 hrs.
===========================
Cheers,
Stan H.